As a financial manager for Micro Electronics, Inc, you have the following information:
i) The company follows a residual dividend policy;
ii) The company has a Capital budget ─ $900,000
iii) The company’s Target capital structure ─ 30% debt, 70% equity;
iv) Forecasted net income ─ $800,000
Please answer the following questions:

a. What will be the amount funded with equity for the project?

b. Compute the amount of the dividend.

As a financial manager for Micro Electronics, Inc, you have the following information:
i) The company follows a residual dividend policy;
ii) The company has a Capital budget ─ $900,000
iii) The company’s Target capital structure ─ 30% debt, 70% equity;
iv) Forecasted net income ─ $800,000
Please answer the following questions:

a. What will be the amount funded with equity for the project?  (sample answer: $250,000)
b. Compute the amount of the dividend. (sample answer: $250,000)
c. Compute the dividend pay-out ratio.
Answer #1

PT a Capital Budget = 900000

Equity Finance =900000*70% i.e. 630000

PT b Income =800000

Dividend =800000-630000 i.e.170000

PT c dividend pay-out ratio =Dividend/earning

=170000/800000 i.e.0.2125 i.e.21.25%